OpenStudy (anonymous):

what is 3/21 in simplest form (ratio)

OpenStudy (kidrah69):

\[\frac{ 3\div3=? }{ 21\div3=? }\]

OpenStudy (anonymous):

3/3=1 21/3=7

OpenStudy (anonymous):

@kidrah69 ?

OpenStudy (kidrah69):

So then just put it into ratio form do you know how?

OpenStudy (anonymous):

no @kidrah69

OpenStudy (kidrah69):

Well ratios have : <---colons right? so if this was a ration it would be 3:21 so how would you put 1/7 as a ratio?

OpenStudy (anonymous):

1:7 @kidrah69

OpenStudy (kidrah69):

Right are there options for this?

OpenStudy (anonymous):

no @kidrah

OpenStudy (kidrah69):

Then i would go with 1:7

OpenStudy (anonymous):

@kidrah69 what is the simplest form of 16/24 (ratio)

OpenStudy (kidrah69):

Repost that